OpenStudy (anonymous):

how do i solve 3+8y=3y-2 i don't think i'm solving it right because my book has a different answer

OpenStudy (anonymous):

\[3 + 8y = 3y - 2\] \[3+5y=-2\] \[5y=-5\] \[y=-1\]

OpenStudy (anonymous):

oh ok i kinda get it, can you send me another problem to just make sure im understanding this right

OpenStudy (anonymous):

Sure, I'm making this one up, but try to work through it, rearranging to isolate the variable on one side. \[2x+5=-x+4\] Post your work, and I'll see if I can point out where you get snagged up...that is if you do get snagged up!

OpenStudy (anonymous):

\[2x+5=-x=4\] \[1x+5=4\] \[1x=-1\] \[x=0?\]

OpenStudy (anonymous):

or is it 1

OpenStudy (anonymous):

Alright, not the correct answer, but I can point right to where you slipped up. Since on the right side, we have a -x+4, which is a -x, we have to add x to both sides to remove it: \[(2x+5)+x = (-x+4)+x\] \[3x+5=4\] See if you can finish it from there.

OpenStudy (anonymous):

\[3x+5=4\] \[3x=-1\] x=-4? because you have to get x by itself so i would minus 3 from both sides?

OpenStudy (anonymous):

Since 3x is = (3*x), we can't just subtract 3 from both sides. In this case, we'll have to divide to undo the multiplication: \[(3x)/3 = -1/3\] \[x = -1/3\] Sorry if you aren't firmilar with fractions...I made this problem up off the top of my head, didn't check fro round numbers! Either way, you'll see them soon enough. Don't let them intimidate you! (They always [and still do] scared me to death.)

OpenStudy (anonymous):

To wrap things up, here the solution straight through: \[2x+5=-x+4\] \[3x+5=4\] \[3x=-1\] \[x=-1/3\] Does that make sense?

OpenStudy (anonymous):

i had -1 over 3 but thought it should have been a whole number

OpenStudy (anonymous):

yeah thanks so much, i guess i should just trust my work even if it does not come out as a whole number

OpenStudy (anonymous):

Perfect! Happy to hear it. Quickly you'll find things are rarely in whole numbers! Best of luck to you!

OpenStudy (anonymous):

thanks again

OpenStudy (anonymous):

:):):):):):) Good Luck